OpenStudy (anonymous):

The real solutions of x^2-4x-2=0. Show all steps please!

OpenStudy (luigi0210):

You have to use the quadratic formula if I'm not mistaken

OpenStudy (anonymous):

Yeah I know that much but I'm not sure how to do it

jimthompson5910 (jim_thompson5910):

I'll get you started \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-4)\pm\sqrt{(-4)^2-4(1)(-2)}}{2(1)}\] \[\Large x = \frac{4\pm\sqrt{16-(-8)}}{2}\]

OpenStudy (luigi0210):

Yea.. What Jim did.. :/

OpenStudy (anonymous):

Thanks so much!!!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."